OpenStudy (anonymous):

Solve by the substitution method. 2x+9y=20 x=8-4y What is the solution of the system, if there is one? A. The solution is ____. B. There is no solution.

OpenStudy (anonymous):

y=-36 and x=-136

OpenStudy (anonymous):

sub x=8-4y into the top equation, 2(8-4y) + 9y =20 16- 8y + 9y =20 y=20-16 y=4 sub y =4 into the 2nd equation, x =8-4(4) x=-8 Therefore, the solution for the system is x =-8, y=4

OpenStudy (anonymous):

Sorry for wrong answer.I didn't notice the sign!!!